### v1

# EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery

Abstract
Recent advances in the visual-language area have developed natural multi-modal large language models (MLLMs) for spatial reasoning through visual prompting. However, due to remote sensing (RS) imagery containing abundant geospatial information that differs from natural images, it is challenging to effectively adapt natural spatial models to the RS domain. Moreover, current RS MLLMs are limited in overly narrow interpretation levels and interaction manner, hindering their applicability in real-world scenarios. To address those challenges, a spatial MLLM named EarthGPT-X is proposed, enabling a comprehensive understanding of multi-source RS imagery, such as optical, synthetic aperture radar (SAR), and infrared. EarthGPT-X offers zoom-in and zoom-out insight, and possesses flexible multi-grained interactive abilities. Moreover, EarthGPT-X unifies two types of critical spatial tasks (i.e., referring and grounding) into a visual prompting framework. To achieve these versatile capabilities, several key strategies are developed. The first is the multi-modal content integration method, which enhances the interplay between images, visual prompts, and text instructions. Subsequently, a cross-domain one-stage fusion training strategy is proposed, utilizing the large language model (LLM) as a unified interface for multi-source multi-task learning. Furthermore, by incorporating a pixel perception module, the referring and grounding tasks are seamlessly unified within a single framework. In addition, the experiments conducted demonstrate the superiority of the proposed EarthGPT-X in multi-grained tasks and its impressive flexibility in multi-modal interaction, revealing significant advancements of MLLM in the RS field.
Index Terms:
Multi-source, Spatial reasoning, Remote sensing (RS), Multi-modal large language models (MLLMs).
I Introduction
Typical spatial understanding tasks in multi-modal learning include referring and grounding, which essentially require aligning regions in the image with text semantics[1]. Referring aims to accurately generate the interpretation sequence of the specific areas. Grounding requires localizing the objects or rendering object segmentation masks based on the user’s language expression[2, 3]. Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focal of research in the natural scenarios[4, 5, 6]. However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[7]. Although current RS MLLMs can handle a wide range of RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception. They solely rely on text instructions which can lead to ambiguity and limited context. It is indispensable for establishing a comprehensive and spatial-aware MLLM to interpret RS images under flexible interactions.
Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[10]. Inspired by this, universal MLLM and multi-grained prompting learning have been extensively studied in the visual-language learning community[11]. Remarkable advances of MLLMs include close-source Open-AI’s GPT series[12, 13, 14], open-source Meta’s LLaMA series[15, 16], and parameter-efficient LLaMA-Adapter series[17, 18], which demonstrate versatile capabilities of handling multi-modal various visual tasks. On top of that, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[19]. Specifically, representative works comprise promptable segmentation model segment anything (SAM)[20], region-of-interest perception model GPT4RoI [4], refer-and-ground MLLM Ferret[1], etc. However, the aforementioned models are tailored for the natural imagery, it is challenging to directly adapt them to the diverse RS data.
Figure 1: An overview of the proposed flexible and comprehensive spatial understanding framework for RS imagery: (a) Task: various multi-grained referring and grounding tasks; (b) Model: the proposed model consists of an image encoder, a visual prompts encoder, a text tokenizer, a LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is transformed based on existing multi-source RS data including optical, SAR, and infrared.
Limited efforts in the RS domain have been made to enhance spatial understanding. A key example is RSPrompter [21], which is based on SAM and develops an automated prompts generation approach to facilitate interactive segmentation. However, RSPrompter is only designed for instance segmentation. RSVG [2] defines the visual grounding tasks in the RS domain and can locate the referred objects using boxes under natural language expression. Recent notable works, such as EarthGPT[8] and Geochat[9], have realized RS holistic and regional imagery comprehension under the unified visual-language learning framework, facilitating the development of MLLMs in the RS domain. Nevertheless, RSVG, EarthGPT, and Geochat solely support language interaction without visual prompts to realize referring expression comprehension. Notably, EarthMarker[22] is constructed as the first visual prompting learning framework in the RS domain and is capable of region-level and point-level RS imagery interpretation. Regretfully, EarthMarker is mainly oriented towards optical images and lacks support for multi-source imagery and free-form visual prompts, hampering flexible human-AI interaction and comprehensive understanding of RS data. Thus, the study on enabling MLLMs to flexibly interpret multi-source RS images at any granularity remains in its infancy stages.
To resolve the limitations of existing models and draw inspiration from how humans understand multi-modal, multi-grained information through flexible interactions, a spatial MLLM named EarthGPT-X is proposed based on visual prompt learning. EarthGPT-X is designed for a comprehensive understanding of multi-source RS imagery under flexible visual prompts (e.g., points, boxes, circles, and scribbles), and unifying referring and grounding tasks within a free-form region-text alignment architecture. As depicted in Fig. 1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, a construction pipeline of training datasets. Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification). Secondly, the key of the model is to realize multi-modal content integration including images, visual prompts, and text instructions. The multi-modal signals are then fed into the large language models (LLM) for subsequent mixed tuning. Additionally, to unify referring and grounding in a single framework, the pixel perception module is designed to generate binary segmentation masks related to referring objects. Thirdly, the training dataset named M-RSVP is constructed, transforming from existing multi-source datasets, which is uniformly reorganized and relabeled as image-visual prompts-text triples, with a portion re-annotated using GPT-4V[23] to support the generation of longer and more detailed descriptions[22].
In addition, observed that existing synthetic aperture radar (SAR) and infrared datasets are limited to object detection tasks, which are insufficient for developing complex visual reasoning capabilities of MLLMs. To address this shortcoming, the multi-domain data are integrated for generalizing general-domain knowledge into SAR and infrared scenarios. Furthermore, to allow for free-form visual prompts data, a noise augmentation method is proposed to add noise that either partially covers the ground truth or extends beyond it to simulate free-form shapes. Based on the newly constructed training datasets, a one-stage all-in-one training strategy is proposed for mixed tuning, facilitating cross-domain data mutual learning and alleviating the need for manual effort for manually setting during the training.
To further investigate EarthGPT-X’s superior comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs. Impressively, EarthGPT-X can realize multi-source data reasoning accurately and provide a more comprehensive and detailed understanding of RS images. Furthermore, to evaluate the effectiveness of the EarthGPT-X’s structure, we conduct ablation studies to assess the designed components’ contributions to the overall performance. Experimental results demonstrate that the proposed EarthGPT-X not only achieves competitive performances on vision-language tasks, but also creates impressive results on referring and grounding tasks.
Our contributions can be summarized as follows.
  • A RS Multi-source Spatial MLLM, EarthGPT-X. A pioneering MLLM named EarthGPT-X is proposed, specifically designed for a comprehensive spatial understanding of multi-source RS imagery. EarthGPT-X advances existing RS MLLMs by enabling flexible multi-grained imagery interpretation, and offers zoom-in and zoom-out insight, powered by designed free-form visual prompts. Furthermore, EarthGPT-X unifies referring and grounding tasks into a visual prompting learning framework, facilitates an adaptable understanding of RS data, and represents a significant milestone in developing versatile MLLMs on RS scenarios.
  • First Multi-source Visual Instruction Dataset, M-RSVP. A multi-source multi-modal visual-prompt-based instruction dataset named M-RSVP, comprising over 200 k image-visual prompts-text triples, is constructed. M-RSVP fills the gap of multi-source visual prompting data in the RS domain, facilitating a multi-grained understanding of RS imagery across various scenes, diverse sources, and varying imaging conditions.
  • Flexible Multi-modal Content Interaction. A multi-modal content integration mechanism is developed, enhancing the mutual understanding between dense image representations, sparse visual prompt features, and text instruction embeddings. This mechanism leads to a deeper comprehension of the referring regions in holistic images. Moreover, by combining with a noise augmentation method for simulating free-form shapes, this mechanism endows the proposed EarthGPT-X with comprehensive RS imagery understanding under flexible human-AI interaction.
  • Cross-domain Knowledge Integration. A cross-domain one-stage fusion training strategy is proposed, transferring and generalizing universal domain-level knowledge into SAR and infrared scenarios. This approach effectively mitigates the limitations of existing SAR and infrared data interpretation, which are typically restricted to object detection tasks and hinder the potential to perform complex multi-task visual reasoning. Furthermore, the one-stage all-in-one training facilitates cross-domain data mutual perception and reduces the reliance on extensive manual effort.
II Related Work
II-A LLMs and MLLMs in Natural Scenarios
The rapid development of LLMs has significantly advanced natural language processing (NLP), demonstrating remarkable capabilities in diverse contexts. Specifically, GPT-3 [12] has demonstrated the power of large-scale models, containing billions to trillions of parameters. InstructGPT [24] and ChatGPT [13] have further enhanced this by showcasing remarkable fluency and adaptability in various conversational tasks. Impressively, the open-source community has significantly contributed to advancing LLMs, with resources like LLaMA series [25] enhancing the instruction-following capabilities of these models. Recent efforts, including Alpaca [26], Vicuna [27], and GPT-4-LLM [28], have introduced full fine-tuning techniques to effectively equip LLMs with robust instruction-following abilities. Additionally, LLaMA-Adapter [18] has demonstrated that parameter-efficient fine-tuning can achieve comparable results with full parameter fine-tuning, offering a more efficient approach to model adaptation.
Notably, an essential step toward achieving general intelligence is integrating multi-modal perception into LLMs, allowing them to handle multi-modal data beyond text. Several works [29, 30, 31, 32] focus on incorporating multi-modal inputs to support task-specific objectives across various modalities. Models like VisualGPT [29], BLIP [30], Flamingo [32], and Kosmos-1 [33] have shown strong multi-modal reasoning abilities by aligning LLMs with visual data. This alignment is achieved using techniques like projection layers, zero-shot attention mechanisms, and intermediate networks, as demonstrated by MiniGPT-4 [34], LLAMA-Adapter V2 [17], and mPLUG-Owl [35]. In addition, with multi-modal LLMs continuing to evolve, other modalities include audio, video [36, 37], and 3D point clouds [38, 39, 40] are being integrated into MLLMs.
II-B MLLMs in RS Scenarios
Inspired by MLLMs in natural domains, MLLMs in the RS field have also made significant strides. Several pioneering RS MLLMs have emerged recently. For example, RSGPT [41], which fine-tunes Instruct-BLIP [42] on a high-quality image-text pairs dataset, demonstrates strong capabilities in image captioning and visual question answer (VQA). However, RSGPT struggles with other visual tasks including object classification and detection, lacking generalization abilities. A key advance is Geochat [9], which is the first versatile MLLM designed to solve multiple tasks for optical RS images. EarthGPT [8] introduces a universal MLLM capable of handling multi-source RS imagery and a broad range of RS visual tasks, and supports regional visual reasoning. In addition, another RS MLLM called LHRS-Bot[43] has established the LHRS-Bench to assist the RS community in evaluating RS-domain MLLMs across diverse evaluation metrics. These models have undoubtedly advanced the development of MLLMs within the RS-specific domain. Nevertheless, EarthGPT, Geochat, and LHRS-Bot solely allow language interaction in multi-turn conversations to handle visual tasks, and fail to support referring comprehension under visual prompt instruction, which restricts flexible interaction in real-world applications.
II-C Prompting Learning
Prompt learning has emerged as a key area of research in NLP [12]. Notable works such as AutoPrompt [44] and CoOp [45] focus on automating prompt template generation for language and vision-language models, reducing the need for manual crafting. In parallel, language prompting has been applied to open-vocabulary detection models, including DetPro [46] and PromptDet [47]. While language prompting is well-developed, visual prompting remains underexplored. A significant advancement in this area is the SAM [20] model, which supports various segmentation prompts to improve zero-shot performance. However, SAM lacks object semantic labels, which is addressed by Semantic-SAM [48]. Moreover, models like GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition. Colorful Prompting Tuning (CPT) [49] uses color-based prompts to enhance MLLMs performance. Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding. Ferret [1] is capable of understanding spatial referring of any shape or granularity within an image and accurately grounding open-vocabulary descriptions. Additionally, SPHINX-V [50] supports free-form visual prompts to achieve pixel-level image comprehension. Mentioned above models are designed for the natural domain.
Note that EarthMarker [22] stands as the pioneering visual prompting framework tailored for the RS domain, specifically designed to interpret RS imagery at both region and point levels. EarthMarker constructs a large-scale RS visual prompting dataset and leverages visual prompts learning to endow MLLM with region-level and point-level RS imagery interpretation capabilities. However, EarthMarker is primarily focused on optical images and lacks support for multi-source imagery and free-form visual prompts. This limitation significantly hinders its potential for flexible interactions and restricts its ability to provide a comprehensive understanding of the diverse and complex characteristics of RS data and broader applicability in real-world RS scenarios.
In summary, existing RS MLLMs have constructed the visual-language learning diagram to comprehend RS imagery and support language interaction in multi-turn dialogues. More advanced than existing MLLMs and EarthMarker, the proposed EarthGPT-X is designed for multi-source imagery comprehension using a dual-prompt system of text instructions and visual prompts. With free-form visual prompting, EarthGPT-X offers comprehensive and flexible visual perception. A straightforward side-by-side comparison of EarthGPT-X versus the recent RS MLLMs is shown in Tab. I.
TABLE I: Comparisons of EarthGPT-X v.s. Recent MLLMs in RS domain.
Model Text Image Visual Prompts Pixel-level
Instruction Optical Multi-source Point Box Free-form Grounding
RSGPT[41]
RS-LLaVA[51]
GeoChat[9]
EarthGPT[8]
Popeye[52]
SkyEyeGPT[53]
EarthMarker[22]
EarthGPT-X
III Methodology
The goal is to unify the referring comprehension at any granularity and pixel grounding into a visual prompt framework in the multi-source RS scenarios. To this end, EarthGPT-X is developed to explore a more general and comprehensive visual prompting learning paradigm. The entire EarthGPT-X framework is detailed in this section. Specifically, we first elaborate on the multi-modal content integration mechanism (Section III-A). Subsequently, the cross-domain one-stage fusion training strategy is presented (Section III-B). In addition, the pixel perception module is introduced in Section III-C.
III-A Multi-modal Content Integration
Figure 2: Overview of the proposed EarthGPT-X model architecture.
The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module. Those multi-modal content encoding and integration are detailed as follows.
III-A1 Visual Global and Local Perception
As shown in Fig. 2 (a), to enhance visual comprehension, the high-resolution inputted images are divided into sub-images for local visual perception. Concurrently, the images are downsampled into lower resolutions and combined with the sub-images for subsequent mixed encoding. The Mixed Visual Encoders (MVE) are designed to encode inputted images and their variants $I^{i}$ into visual representations $V_{\mathrm{img}}$. To leverage the advantages of different network architectures (CNN and Vision Transformer (ViT) [54]) and integrate the local information and contextual dependencies, the semantically complementary backbones CLIP-ConvNeXt [55] and ViT self-supervised by DINOv2[56] are employed as MVE. This visual encoding process can be formulated as
$V_{\mathrm{img}}=\mathrm{Concat}~{}(\mathrm{MVE}~{}(I^{i})),~{}i=1,2,...,N.$ (1)
III-A2 Free-form Spatial-aware
Free-form visual prompts are expected to be flexibly used to guide the model in generating appropriate responses. The shapes of visual prompts include points, boxes, circles, scribbles, etc. To meet the need for flexible interaction at any granularity, a noise augmentation method is developed. Specifically, Gaussian distribution based on the dimensions of the box prompts is sampled to simulate the free-formed shapes. The introduced noise encompasses only a portion of the ground truth or expands it. Subsequently, considering that the visual prompts exhibit varying degrees of sparsity across different shapes, as illustrated in Fig. 2 (b), a randomly initialized ViT is employed as a spatial-aware encoder (SAE) specifically for feature extraction, without any pre-trained weights suitable for image features. The various types of visual prompts $V_{\mathrm{p}}$ are encoded using the designed SAE. The implementation equation can be expressed as follows
$E_{\mathrm{prompt}}=\mathrm{SAE}~{}(V_{\mathrm{p}}),$ (2)
where $E_{\mathrm{prompt}}$ denotes the encoded visual prompts features with free-form shapes.
III-A3 Text Instructions Tokenizer
As displayed in Fig. 2 (c), To facilitate linguistic comprehension, the text instructions $T_{\mathrm{orig}}$ are encoded by the tokenizer[57] module, and projected into language embeddings $L_{\mathrm{instruct}}$, preparing for the subsequent LLM to capture the nuanced interdependence. The detailed process after disassembly is first to use the subword tokenization[57] method to divide text instructions into subwords, enabling understanding of the structure of the linguistic signals. Subsequently, the Word2Vec[58] is adopted to map each subword to a one-dimensional vector embedding. The tokenizer process can be simply described as
$L_{\mathrm{instruct}}=\mathrm{Tokenizer}(~{}T_{\mathrm{orig}}).$ (3)
III-A4 Hybrid Signals Mutual Understanding
To bridge the gap and enhance the interplay between dense visual representations and sparse visual prompts features, a hybrid signal mutual understanding mechanism is proposed. Specifically, as displayed in Fig. 2 (d), the image representations $V_{\mathrm{img}}$ and visual prompts features $E_{\mathrm{prompt}}$ are processed by a self-attention layer and feed-forward layer to obtain $V_{\mathrm{sa}}$ and $E_{\mathrm{ff}}$, respectively. The subsequent cross-attention and second self-attention blocks are used to compute the correlation $H_{\mathrm{ca}}$ between hybrid referring information and multi-scale holistic signals. In essence, this designed mechanism enhances detailed and contextual visual perception, leading to a deeper comprehension of the referring areas in holistic images. The whole procedure can be mathematically represented as
$V_{\mathrm{sa}}$ $=\mathrm{Self\text{-}attention}~{}(V_{\mathrm{img}}),$ (4)
$E_{\mathrm{ff}}$ $=\mathrm{Feed\text{-}forward}~{}(E_{\mathrm{prompt}}),$ (5)
$H_{\mathrm{ca}}$ $=\mathrm{Cross\text{-}attention}~{}(V_{\mathrm{sa}},~{}E_{\mathrm{ff}}),$ (6)
$[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~{}(V_{\mathrm{ca}}+R_{\mathrm{ff}}),$ (7)
where $[V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts. Then, the hybrid visual features are projected to the same dimension as text embeddings through the vision-to-language modality-align projection layer $\mathcal{F}_{\mathrm{vl\text{-}p}}$. This process can be formulated as
$[V\text{-}P]_{\mathrm{proj}}=\mathcal{F}_{\mathrm{vl\text{-}p}}~{}([V\text{-}P]_{\mathrm{hybrid}}).$ (8)
where $[V\text{-}P]_{\mathrm{proj}}$ involve both multi-scale visual signals and free-form spatial-aware information. Notably, the cross-attention layer is updatable during the training.
III-B Cross-domain One-Stage Fusion Training
The previous training pipeline of MLLM routinely uses a step-wise training strategy[8, 22]. However, in the previous pipeline, extensive manual intervention is required. For example, the researcher needs to manually set the different trainable parameters and datasets across various stages. To alleviate the need for manual efforts without sacrificing performance, the cross-domain one-stage fusion training[11] strategy is designed. Most importantly, this strategy further facilitates multi-domain (i.e., natural, optical, SAR, and infrared) data mutual understanding.
In this stage, the projected hybrid tokens are concatenated with the text instruction embeddings $L_{\mathrm{instruct}}$ to develop the multi-modal LLM input $\mathcal{X}$. Subsequently, the integrated multi-modal input is injected into LLM for alignment and deeply interactive understanding (see Fig. 2 (e)). The process can be expressed as
$Y=\mathrm{LLM}~{}(~{}[V\text{-}P]_{\mathrm{proj}},~{}L_{\mathrm{instruct}}).$ (9)
Unlike other models that introduce additional learnable parameters, to simplify the model structure, the Transformer’s attention layers are unfrozen during the training. Specifically, the self-attention head consists of key $K$, query $Q$, and the value $V$, which are implemented by linear layers. The implementations can be expressed as follows
$Q(\mathcal{X})=W_{\mathrm{q}}~{}\mathcal{X}+{b_{\mathrm{q}}},$ (10)
$K(\mathcal{X})=W_{\mathrm{k}}~{}\mathcal{X}+{b_{\mathrm{k}}},$ (11)
$V(\mathcal{X})=W_{\mathrm{v}}~{}\mathcal{X}+{b_{\mathrm{v}}},$ (12)
where $\mathcal{X}$ denotes multi-modal input. The parameters $W_{\mathrm{q}}$, $W_{\mathrm{k}}$, $W_{\mathrm{v}}$, $b_{\mathrm{q}}$, $b_{\mathrm{k}}$, and $b_{\mathrm{v}}$ are updatable during the training.
To summarize, the essence of this process lies in levering multi-domain knowledge to generalize versatile visual reasoning ability into RS multi-source complex scenarios interpretation. Furthermore, through tuning on unified abundant visual prompt-based instruction data, the proposed EarthGPT-X can accomplish referring comprehension of RS imagery with the user’s free-form visual prompts and text instruction-following.
III-C Pixel Perception
To unify referring and grounding in one framework and achieve pixel-level grounding, the pixel encoder-decoder architecture is employed to render the referring object masks, which are shown in Fig. 2 (f). In particular, the special token $pix$ is adopted to represent the pixel grounding masks as embeddings[59]. The embeddings $L_{\mathrm{pix}}$ are generated from the LLM output corresponding to the $pix$ token. Then, the language-to-pixel projection layer $\mathcal{R}_{\mathrm{lp\text{-}p}}$ transforms the embeddings $L_{\mathrm{pix}}$ into the pixel decoder’s feature space. In addition, the grounding encoder is denoted as $\mathcal{G}$, the pixel decoder $\mathcal{P}$. The binary segmentation mask $M$ generated process can be represented as follows
$M=\mathcal{P}\Big{(}\mathcal{R}_{\mathrm{lp}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big{)}.$ (13)
In summary, based on the designed tripartite architecture and all-in-one training strategy, the proposed EarthGPT-X can be expected to effectively and comprehensively realize cross-modal and multi-source knowledge comprehension. Moreover, EarthGPT-X can handle multi-grained RS visual tasks, and flexibly unifies referring and grounding in one framework.
IV Dataset Construction
The goal is to endow the proposed model with flexible and comprehension interpretation abilities on multi-source imagery. However, existing infrared and SAR datasets are limited to object detection tasks, hindering the development of various advanced reasoning capabilities. To address this limitation and unleash the potential of MLLMs to comprehend diverse RS data, a multi-source visual prompt-based instruction dataset named M-RSVP is constructed. Particularly, the natural and optical RS image data, which involves a wide range of visual tasks, are utilized for LLM mixed tuning. In essence, multi-domain data fusion facilitates generalizing the universal domain reasoning capabilities into limited SAR and infrared scenarios. The specific M-RSVP construction pipeline is detailed as follows.
IV-A Raw Data Selection
In order to develop multi-source image understanding abilities of the proposed EarthGPT-X, it is indispensable to utilize data from different scenes and sensors, including outdoor natural images, RS optical images, SAR, and infrared images, covering different visual task types. Thus, the natural scene dataset we selected is OpenImages[60] which meets this characteristic and is ideal for training the model’s visual perception capabilities in general environments. Moreover, we also choose Visual Genome[61] for training, to practice reasoning over multi-region and multi-object interactions. Meanwhile, for the RS optical dataset, we use RSVP[22] which is the first RS visual prompting dataset, allowing the model to develop region/point interpretation skills in diverse geographic and environmental settings. Furthermore, SAR detection datasets includes SSDD[62], AIR-SARShip-2.0 [63], SARDet [64], MSAR [65], and SAR-Aircraft [66]. Infrared object detection contains Sea-shipping, Infrared-security [67], Aerial-mancar [68], Double-light-vehicle [69], and HIT-UAV [70]. Leveraging multi-domain datasets is beneficial for equipping the proposed model to obtain robust generalization capabilities and to handle a wide range of visual tasks. In addition, multi-source data integration ensures comprehensive and flexible image understanding across multiple domains and sensor types.
IV-B Multi-source Data Unification
[…]
In conclusion, all the various datasets from different domains are combined for mixed training to deepen mutual understanding. The multiple free-form visual prompts are allowed to instruct the model to generate suitable answers, ensuring the flexibility of human-AI interaction. Furthermore, multi-domain data fusion is beneficial to realize multi-source multi-task interpretation by leveraging the diverse features and scenarios presented in supplementary image types.
V Experiment
V-A Experimental Setups
V-A1 Implementation Details
The one-stage all-in-one training is conducted in this paper. We use transformer-based decoder-only Llama 2-13B[16] as the basic LLM. The pre-trained DINOv2-ViT L/14 [56] and CLIP-ConvNeXt [55] are adopted as hybrid image encoder MVE. The training image resolution is $448\times 448$. Subsequently, the randomly initialized ViT model is seated as the spatial-aware encoder to refine sparse features. Then, we employ a pre-trained SAM encoder and decoder[20] as the grounding encoder and pixel decoder. The image resolution required to be reshaped for the segmentation model is $1024\times 1024$. Furthermore, we utilize the simple MLPs as the projection layer to realize multi-modal content dimension alignment.
Throughout the training, to keep training efficiency and preserve the generalization ability of backbone models, we only unfreeze the self-attention layers of LLM and the added cross-attention module for tuning. Other modules including visual encoders, projection layers, etc., are kept frozen. The AdamW optimizer[72] with a weight decay of 0.01 and a learning rate of 2e-5 is utilized. The total training is conducted on 8 NVIDIA A100 GPUs with 80 GB memory.
V-A2 Evaluation Metrics
We adopt Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS)[73] as evaluation metrics on referring object classification tasks. Furthermore, we evaluate the performance of the proposed model on the region captioning tasks based on metrics for assessing language generation quality, including BLEU[74], METEOR[75], ROUGE[76], CIDER[77], and SPICE[78]. Those metrics provide a diverse and comprehensive assessment of the quality of identified object information and generated region captions.
TABLE II: Results on referring object classification of EarthGPT-X and other multiple MLLMs on DIOR-RSVG test set.
Methods Publication Year Formats SS S-IoU
Vanilla MLLMs
GeoChat [9] CVPR 2024 Coor 79.59 68.80
Sphinx [79] Arxiv 2023 Coor 93.72 89.37
EarthGPT [8] TGRS 2024 Coor 94.64 90.16
Natural Visual
Prompting MLLMs
Sphinx-V [50] Arxiv 2024 Box 89.07 81.62
Vip-LLava-7b [80] CVPR 2024 Box 72.56 55.94
Vip-LLava-13b [80] CVPR 2024 Box 74.51 60.53
RS Visual
Prompting MLLM
EarthMarker [22] Arxiv 2024 Point 95.96 93.49
EarthMarker [22] Arxiv 2024 Box 98.37 97.24
EarthGPT-X(Ours) Point 96.02 94.12
EarthGPT-X(Ours) Box 98.60 97.83
EarthGPT-X(Ours) Free-form 98.30 96.56
TABLE III: Results on region captioning of EarthGPT-X and other MLLMs on the OPT-RSVG test set.
Method Publication Year Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDER SPICE
Vanilla MLLMs
Qwen-VL-Chat [9] Arxiv 2023 Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
GeoChat [81] CVPR 2024 Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
Sphinx [79] Arxiv 2023 Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
EarthGPT [8] TGRS 2024 Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
Visual Prompting MLLMs
Sphinx-V [50] Arxiv 2024 Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
Vip-LLava-7b [80] CVPR 2024 Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
Vip-LLava-13b [80] CVPR 2024 Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
GLaMM [80] CVPR 2024 Box 12.50 6.06 3.25 1.74 7.98 22.01 23.85 12.40
EarthGPT-X(Ours) Point 52.22 48.13 45.15 42.88 35.01 64.31 477.40 63.14
EarthGPT-X(Ours) Box 64.55 59.76 56.10 53.11 39.22 66.64 487.44 64.20
EarthGPT-X(Ours) Free-form 63.25 58.16 54.89 52.01 38.35 65.01 482.26 63.23
Figure 3: The referring object classification results on multi-source RS images demonstrate the superior region-level RS visual understanding capability of EarthGPT-X, compared to other vanilla and visual prompting MLLMs (the black text indicates consistency with the ground truth, the red text represents incorrect answers).
Figure 4: Example visualization of EarthGPT-X for multi-source (optical, SAR, and infrared) image comprehension and multiple visual tasks reasoning under various visual prompts.
V-B Referring Object Classification
To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla MLLMs and visual prompting MLLMs on DIOR-RSVG[7]. For referring object classification tasks, the text instruction is “Please identify the object category of the marked region in the image”. The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions. The vanilla MLLMs including natural MLLMs Sphinx [79], RS MLLMs GeoChat [81], and EarthGPT [8]. We select the most recent natural visual prompting MLLMs including Sphinx-V [50], Vip-LLava [80], and the RS visual prompting MLLM model EarthMarker[22] as the comparison.
[…]
Figure 5: Pixel grounding examples of EarthGPT-X on multi-source imagery (EarthGPT-X’s interpretation of the object classes is also displayed on the figures).
V-C Region Captioning
For the brief region captioning, the test set of OPT-RSVG [82] is employed for zero-shot evaluation. We elaborately choose both natural and RS-specific MLLMs for performance comparison. Natural scene MLLMs contains Qwen-VL-Chat [9], Sphinx [79], Sphinx-V [50], Vip-LLava [80] and GLaMM [80]. Popular RS-specific MLLMs GeoChat [81] and EarthGPT [8] are selected. The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”. As displayed in Tab. III, the proposed EarthGPT-X significantly outperforms the previous works on all of the evaluation metrics. Notably, using boxes as visual prompts can yield the best performance on region captioning and dialogue ability.
V-D Comprehensive Capabilities Evaluation
Traditional vision models in the RS field primarily focus on a single visual modality and single tasks, exhibiting poor generalization[83, 84, 85]. Thanks to our M-RSVP’s rich multi-sensor knowledge, EarthGPT-X has robust generalization capabilities across optical, SAR, and infrared visual modalities. To verify and assess the comprehensive capabilities of the proposed model, we conduct a thorough qualitative evaluation. Our assessment also covers multi-source images and multiple levels of image interpretation granularity, including image, point, box, and free-form prompted levels. Moreover, the evaluated visual tasks are extensive, such as brief captioning, detailed captioning, ground captioning, region captioning, referring object classification, and inter-relationship analyses. As displayed in Fig. 4, our model demonstrates remarkable visual reasoning and analytical abilities. EarthGPT-X can provide detailed descriptions for multi-source images, as shown in the interpretation results of the first image, which includes overall image depiction, number of targets, target orientation, and scene reasoning. Notably, in the relationship analysis of the second and third images, EarthGPT-X provides detailed and comprehensive responses. Specifically, EarthGPT-X adeptly summarizes the connections between various regions and the overall image relationship, and then analyzes the characteristics and attributes of the targets within each specific area. In summary, EarthGPT-X demonstrates a profound ability to dissect and understand complex visual information in the multi-source RS MLLMs.
V-E Pixel Grounding
To evaluate the pixel grounding capabilities of the proposed model, we specifically chose images that feature complex and different scenes with multiple targets. This experiment aims to test the model’s robustness and precision in identifying and segmenting diverse elements within a single frame. As shown in Fig. 5, the first test image is the port scene, we mark a ship and a tiny vehicle to assess the model’s ability to discern and accurately interpret tiny targets amidst a busy background. Surprisingly, EarthGPT-X not only recognized these smaller objects but also effectively segmented them from the surrounding environment. Another test involved an image from an urban sports center where we marked five distinct targets scattered across different areas, including varied elements such as a swimming pool, tennis court, vehicle, etc. This scenario tests the model’s capacity to handle spatially separated objects within a complex urban setting. EarthGPT-X demonstrates exceptional performances in precise object recognition and pixel-grounding. Moreover, these tests highlight EarthGPT-X’s advanced pixel understanding capabilities and its potential to enhance RS image analyses across different visual modalities.
V-F Ablation Study
In this section, we analyze the impact of key components of EarthGPT-X in detail.
TABLE IV: Ablation results on different visual encoders.
CLIP CLIP DINOv2 SS S-IoU
ConvNeXt-L RN50$\times$16 ViT-L/14
[…]
TABLE V: Ablation results on Image Pre-processing.
Settings Visual Encoder SS S-IoU
Single-scale MVE 94.93 95.05
Multi-scale MVE 96.18 95.29
Multi-scale + Sub-images MVE 98.40 97.13
V-F1 Different Visual Encoders
To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders. Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison. The results are shown in Tab. IV, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieve the most superior performance. The performance of both SS and S-IoU increases accordingly. These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
V-F2 Sub-images Decomposition and Multi-scale
In the visual feature extraction phase of our approach, we enhance the integration of visual features by incorporating multiple scales and high-resolution sub-images, enabling a more profound comprehension of RS images. To evaluate the effectiveness of our modified feature extraction strategy, we conduct ablation studies focusing on the impact of multi-scale features and high-resolution sub-images on the performance of our EarthGPT-X model. As shown in Tab. V, utilizing the multi-scale and sub-images features setting achieves superior performance over single-scale features. These results demonstrate that the detailed analysis of these high-resolution sub-images and more scales allows for capturing finer details that are crucial for accurate interpretation and analysis.
VI Conclusion
In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources. As the first RS spatial model to unify referring and grounding tasks within a visual prompting framework, EarthGPT-X represents a significant advancement in the application of MLLMs to the RS domain. Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery. Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios, such as SAR and infrared. Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-grained visual understanding at the image, region, pixel levels, etc., allowing for precise and intelligent analyses in real-world applications. In the future, we plan to enrich the SAR and infrared datasets to encompass a broader range of tasks, addressing the current limitation of narrow scope. Additionally, we aim to enable EarthGPT-X to perform more complex visual reasoning tasks across multiple visual modalities, further solidifying EarthGPT-X as a versatile and powerful tool for comprehensive RS imagery analyses.